# An Empirical Study of Catastrophic Forgetting in Large Language Models During Continual Fine-tuning

### v3

@@ Abstract
Catastrophic forgetting (CF) is a phenomenon that occurs in machine learning when a model forgets previously learned information while acquiring new knowledge. As large language models (LLMs) have demonstrated remarkable performance, it is intriguing to investigate whether CF exists during the continual instruction tuning of LLMs. This study empirically evaluates the forgetting phenomenon in LLMs’ knowledge during continual instruction tuning from the perspectives of domain knowledge, reasoning, and reading comprehension. The experiments reveal that catastrophic forgetting is generally observed in LLMs ranging from 1b to 7b parameters. Moreover, as the model scale increases, the severity of forgetting intensifies. Comparing the decoder-only model BLOOMZ with the encoder-decoder model mT0, BLOOMZ exhibits less forgetting and retains more knowledge. Interestingly, we also observe that LLMs can mitigate language biases, such as gender bias, during continual fine-tuning. Furthermore, our findings indicate that ALPACA maintains more knowledge and capacity compared to LLAMA during continual fine-tuning, suggesting that general instruction tuning can help alleviate the forgetting phenomenon in LLMs during subsequent fine-tuning processes.
@@ 1 Introduction
    What types of general knowledge are saliently forgotten?
@@
To address these questions, we conduct an empirical study on various LLLMs, such BLOOMZ, mT0 (Muennighoff et al., 2022), LLAMA Touvron et al. (2023), and ALPACA Taori et al. (2023) to analyze the catastrophic forgetting (CF) problem during continual instruction tuning. We continually train the original LLMs with five instruction tasks and evaluate the retention of the general knowledge within the model from three perspectives: domain knowledge, reasoning, and reading comprehension. Furthermore, we investigate the evolution of bias in LLMs throughout the tuning process. To gain insights into the effect of model architecture, we compare the performance of BLOOMZ with that of mT0 (Muennighoff et al., 2022) (an encoder-decoder model), which is fine-tuned using similar datasets. We also investigate the impact of general instruction tuning on the CF problem by comparing the performance of the initial model with the instruction-tuned version such as (BLOOM, BLOOMZ) and (LLAMA (Touvron et al., 2023), ALPACA (Taori et al., 2023)).
@@
Table 1: Fine-tuning examples in an instruction form.
Our findings reveal that the forgetting problem is generally present in LLMs. Interestingly, as the model scale increases from 1b to 7b parameters, the severity of forgetting intensifies. One potential explanation for this phenomenon is that larger language models exhibit stronger initial performance and, consequently, experience more pronounced performance degradation during continual instruction tuning. Additionally, we observe that the bias in LLMs is mitigated throughout the continual instruction tuning process. When comparing BLOOMZ with mT0 at comparable model scale, we find that BLOOMZ experiences a relatively milder forgetting problem, suggesting that the decoder-only architecture may be better at retaining information during continual instruction tuning. Lastly, empirical results on LLAMA and its instruction-tuned version (i.e., ALPACA) indicate that diverse instruction tuning can help alleviate the CF phenomenon for LLMs in further continual fine-tuning.
@@
    We provide an initial research evidence that the CF problem generally exists in the continual instruction tuning process for different models such as BLOOMZ, mT0, LLAMA and ALPACA. We also show that the model architecture, model scale and general instruction tuning have different effects on the CF problem.
@@ 2 Related Work
Instruction Tuning. Instruction tuning has proven to be effective in aligning responses from pre-trained language models with human intents or preferences (Ouyang et al., 2022; Stiennon et al., 2020; Min et al., 2021). This technique refines a model’s ability to predict a specific response to a given prompt, which may optionally include an instruction that outlines a task for the model. Examples of such models include T0 (Sanh et al., 2021), mT0 (Muennighoff et al., 2022), and BLOOMZ (Muennighoff et al., 2022). It has been demonstrated that instruction tuning can enhance the ability of language models to generalize to unseen tasks without prior exposure (Wei et al., 2021; Sanh et al., 2021). In this work, we focus on fine-tuning LLMs in a continual manner and analyze the catastrophic forgetting (CF) phenomenon during training. Specifically, instructions for a particular type of task (such as generating headlines) are used to tune the LLMs in each training phase, and the model does not have access to previously learned tasks.
Evaluation of CF in Continual Learning. Various training strategies have been proposed to address the problem of catastrophic forgetting (CF) in continual learning (Riemer et al., 2019; Buzzega et al., 2020; Ke et al., 2022; Chen et al., 2022; Luo et al., 2023a). Previous studies have primarily measured CF by evaluating the performance decrease in previously learned tasks during continual learning or the average performance of learned tasks at the end of training. However, Davari et al. (2022) discovered that even when the model performance on previously learned tasks is preserved, the representations still suffer from significant drift due to parameter updates. As a result, they propose using an optimal linear classifier of learned tasks to measure performance, with changes considered as a surrogate to quantify CF. Similarly, Wu et al. (2022) employs layer-wise and task-wise probing to analyze CF in each layer for previously learned tasks. Luo et al. (2023b) propose using a series of probing tasks to evaluate the knowledge stored in LLMs and analyze the generality of the models. However, their study is limited to classification tasks and encoder-only model architectures. To the best of our knowledge, we are the first to evaluate the forgetting of general knowledge in generative large language models during continual instruction tuning.
@@ 3.1 Continual Tasks
Following Scialom et al. (2022), we adopt instruction tasks dissimilar to the training and evaluation tasks of BLOOMZ and mT0. Specifically, we select 5 tasks from Scialom et al. (2022) as follows:
@@
During the instruction tuning, we first add a general prompt template to the beginning of the data: ‘Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request…’ followed by a specific prompt for each task. We adopt the specific prompts designed by Scialom et al. (2022) and 100,000 data samples are used for training. The details of the instruction are shown in Appendix A. For simplicity, we train the model on one instruction task order for the empirical study: Simp $\to$ Emdg $\to$ InqQG $\to$ Exp $\to$ HGen.
@@ 3.2 Evaluation Tasks
Set Elements
DK STEM, Social, Human, Other
Rs BoolQ, PIQA, Winogrande, Hellaswag, MathQA, Mutual
RC RACE-high, RACE-middle
Socioeconomic, Disability, Age
@@ 3.3 Evaluation Metric
The average decrease of $R^{e}_{m}$ is adopted as a surrogate metric to evaluate the forgetting:
$FG_{i}=\frac{1}{|E_{i}|}\sum_{e\in E_{i}}\frac{1}{N}\sum_{m=1}^{N}\frac{R^{e}_{o}-R^{e}_{m}}{R^{e}_{o}}*100\%,$ (1)
@@
For most multi-choice evaluation elements, we adopt the accuracy in the zero-shot setting to measure the model performance, including MathQA, Hellaswag, BoolQ, PIQA, Mutual, Winograde, and RACE. For MMLU, we adopt the 5-shot setting for evaluation. For CrowsPairs, we follow Nangia et al. (2020) to measure the model preference for the stereotypical sentence based on the perplexity of the given stereotypical and anti-stereotypical sentences, where a larger value means a stronger bias in the language.
@@ 4.1 Large Language Models
BLOOMZ (Muennighoff et al., 2022) is a decoder-only model fine-tuned on multilingual tasks with English prompts based on the model BLOOM. The model scales range from 560M to 176b, providing a test bed for analyzing the forgetting phenomenon across different scales. Specifically, in this study, we continually train BLOOMZ on the scale 1.1b, 1.7b, 3b, and 7.1b.  [^4] https://huggingface.co/bigscience/bloomz
@@ 4.2 Implementation
We train our model on 8 GPU (Tesla A100 40G) using the Adam optimizer Kingma & Ba (2014) (the models in 1b level are trained on 4 GPU for saving resources). For all the models, the batch size is 4 on each device, the learning rate is 2e-5, and the scheduler is set constant for BLOOMZ and mT0 following Muennighoff et al. (2022). The max sequence length of the inputs is 512. We train our model 3 epochs and the final checkpoints are used for evaluation.
@@ 5 Experimental Results and Analysis
5.1 Main Results
Figure 3: The model performance of BLOOMZ-7.1b on the instruction task after continually training the model on the corresponding instruction data. We adopt the evaluation metrics following Scialom et al. (2022), which are also shown in the x-axis. BS refers to BLUEScore and R1 refers to Rouge-1.
Firstly, Figure 3 shows the performance of the continually trained BLOOMZ-7.1b model on the instruction tasks compared to the initial model’s performance. For example, after being trained on the Simp training data, the BLOOMZ model achieves a SARI score of 47.24 on the Simp test set, which is 1.98 higher than the initial model’s score. Subsequently, when we continually train the model on the Emdg dataset, the obtained model achieves a BLEU score of 53.3, surpassing the initial model’s performance by 3.62. These improvements demonstrate that the model can benefit from the instruction tuning process and achieve significantly better performance on the instruction tasks. Additional results for other LLMs are presented in Appendix B.
Next, Figure 2 displays the FG values of BLOOMZ-1.1b and BLOOMZ-7.1B. As we can observe, the performance gradually decreases as we continually tune the model with instruction tasks. For instance, the performance of BLOOMZ-7.1b on MMLU-SocialScience in Figure 2 drops from 36.18% to 26.06% after continual training. The declining performance in LLMs indicates the presence of the catastrophic forgetting (CF) problem during the continual instruction tuning process. Moreover, as more instruction tasks are introduced, the general knowledge suffers more significant forgetting. We can also notice that the performance of the BLOOMZ-7.1b model drops more drastically in these evaluation tasks, as we adopt the same y-axis scale for both BLOOMZ-1.1b and BLOOMZ-7.1b in Figure 2. For example, the performance on MMLU-Other drops from 36.18% to 26.35% in BLOOMZ-7.1b, while it drops from 30.58% to 25.97% in BLOOMZ-1.1b after continual training. Although the initial performance of BLOOMZ-7.1b is more substantial compared to that of BLOOMZ-1.1b, the performance of both models ends at relatively similar values, which would result in a more significant FG value for BLOOMZ-7.1B. [^8] Detailed results of some other LLMs during continual instruction tuning are presented in the Appendix C.
@@ 5.1 Main Results
$R^{e}_{o}$ $R^{e}_{-1}$ $FG$
@@
Table 4: FG values for Bias in LLMs during continual instruction tuning. $R^{s}_{o}$ and $R^{s}_{-1}$ refer to the evaluation results at the beginning and the end of instruction tuning.
5.2 Effect of Scales
We visualize the FG values of domain knowledge, reasoning, and reading comprehension with respect to model scales in Figure 4. We can observe that the forgetting phenomenon becomes increasingly severe as the model scale increases. For example, the FG values in domain knowledge are 9.54%, 10.72%, 14.63%, and 18.37% in BLOOMZ-1.1b, 1.7b, 3b, and 7.1b, respectively. BLOOMZ-7.1b suffers the most drastic forgetting. As shown in Table 1, the initial performance $R_{o}^{s}$ is boosted by the increasing model scale, but the final performance is relatively similar across different scales, which may explain the varying extent of forgetting. The same pattern can also be observed in the mT0-1.2b and mT0-3.7b models, as shown in Table 3. Regarding the bias in LLMs, FG values do not correlate with the model scales, which is also reflected in the initial model performance. In other words, there is no evident correlation between the initial performance $R_{o}^{s}$ in bias and the model scales. This finding suggests that the degree of bias in LLMs is not directly related to their size, and increasing the model scale does not necessarily lead to a corresponding increase or decrease in bias.
Figure 5: The comparison of FG between BLOOMZ and mT0 in the comparable model scale.
@@ 5.3 Effect of Model Architecture
We also compare the forgetting phenomenon of different model architectures in Figure 5. As observed, at comparable model scale, BLOOMZ-1.1b and mT0-1.2b achieve similar FG values in domain knowledge, reasoning, and reading comprehension. However, as the scale increases to 3b, BLOOMZ-3b suffers less forgetting compared to mT0-3.7B. For example, the FG value of BLOOMZ-3b is 11.09 which is 5.64 lower than that of mT0-3.7b. These results suggest that BLOOMZ, which has a decoder-only model architecture, can maintain more knowledge during continual instruction tuning. This difference may be attributed to the autoregressive nature of the model or the differences in training objectives. Furthermore, the results imply that as the model scale increases, decoder-only models may suffer from less catastrophic forgetting compared to encoder-decoder models.

### v4

@@ Abstract
Catastrophic forgetting (CF) is a phenomenon that occurs in machine learning when a model forgets previously learned information while acquiring new knowledge for achieving a satisfactory performance in downstream tasks. As large language models (LLMs) have demonstrated remarkable performance, it is intriguing to investigate whether CF exists during the continual instruction tuning of LLMs. This study empirically evaluates the forgetting phenomenon in LLMs’ knowledge during continual instruction tuning from the perspectives of domain knowledge, reasoning, and reading comprehension. The experiments reveal that catastrophic forgetting is generally observed in LLMs ranging from 1b to 7b parameters. Surprisingly, as the model scale increases, the severity of forgetting intensifies in such a model sale range which may result from the much significant initial performance in the larger LLM. Comparing the decoder-only model BLOOMZ with the encoder-decoder model mT0, BLOOMZ exhibits less forgetting and retains more knowledge. Interestingly, we also observe that LLMs can mitigate language biases, such as gender bias, during continual fine-tuning. Furthermore, our findings indicate that general instruction tuning can help alleviate the forgetting phenomenon in LLMs during subsequent fine-tuning.
@@ 1 Introduction
    How to mitigate such forgetting phenomenon?
To address these questions, we conduct an empirical study on various LLLMs, such BLOOMZ, mT0 (Muennighoff et al., 2022), LLAMA (Touvron et al., 2023), and ALPACA (Taori et al., 2023) to analyze the catastrophic forgetting (CF) problem during continual instruction tuning. We continually train the original LLMs with five instruction tasks and evaluate the retention of the general knowledge within the model from three perspectives: domain knowledge, reasoning, and reading comprehension. Furthermore, we investigate the evolution of bias in LLMs throughout the tuning process. To gain insights into the effect of model architecture, we compare the performance of BLOOMZ with that of mT0 (Muennighoff et al., 2022) (an encoder-decoder model), which is fine-tuned using similar datasets. We also investigate the impact of general instruction tuning on the CF problem by comparing the performance of the initial model with the instruction-tuned version such as (BLOOM, BLOOMZ) and (LLAMA (Touvron et al., 2023), ALPACA (Taori et al., 2023)).
Table 1: Fine-tuning examples in an instruction form.
@@
Output: What is the del operator in the Schrodinger equation?
Our findings reveal that the forgetting problem is generally present in LLMs. Surprisingly, as the model scale increases from 1b to 7b parameters, the severity of forgetting intensifies. One potential explanation for this phenomenon is that larger language models exhibit stronger initial performance and, consequently, experience more pronounced performance degradation because of the fitting on the new task during continual instruction tuning. Additionally, we observe that the bias in LLMs is mitigated throughout the continual instruction tuning process. When comparing BLOOMZ with mT0 at comparable model scale, we find that BLOOMZ experiences a relatively milder forgetting problem, suggesting that the decoder-only architecture may be better at retaining information during continual instruction tuning. Lastly, empirical results on LLAMA and its instruction-tuned version (i.e., ALPACA) indicate that diverse instruction tuning can help alleviate the CF phenomenon for LLMs in further continual fine-tuning.
@@
    We provide an initial research evidence that the CF problem generally exists in the continual instruction tuning process for different models such as BLOOMZ, mT0, LLAMA and ALPACA. We also show that the model architecture, and model scale have different effects on the CF problem.
  3. 3.
    Experimental results further show that the general instruction data can help mitigate the CF problem to some extent by experiments.
@@ 2 Related Work
2.1 Instruction Tuning
Instruction tuning has proven to be effective in aligning responses from pre-trained language models with human intents or preferences (Ouyang et al., 2022; Stiennon et al., 2020; Min et al., 2021). This technique refines a model’s ability to predict a specific response to a given prompt, which may optionally include an instruction that outlines a task for the model. Examples of such models include T0 (Sanh et al., 2021), mT0 (Muennighoff et al., 2022), and BLOOMZ (Muennighoff et al., 2022). It has been demonstrated that instruction tuning can enhance the ability of language models to generalize to unseen tasks without prior exposure (Wei et al., 2021; Sanh et al., 2021). In this work, we focus on fine-tuning LLMs in a continual manner and analyze the catastrophic forgetting (CF) phenomenon during training. Specifically, instructions for a particular type of task (such as generating headlines) are used to tune the LLMs in each training phase, and the model does not have access to previously learned tasks.
2.2 Evaluation of CF in Continual Learning
Various training strategies have been proposed to address the problem of catastrophic forgetting (CF) in continual learning (Riemer et al., 2019; Buzzega et al., 2020; Ke et al., 2022; Chen et al., 2022; Luo et al., 2023a). Previous studies have primarily measured CF by evaluating the performance decrease in previously learned tasks during continual learning or the average performance of learned tasks at the end of training. However, Davari et al. (2022) discovered that even when the model performance on previously learned tasks is preserved, the representations still suffer from significant drift due to parameter updates. As a result, they propose using an optimal linear classifier of learned tasks to measure performance, with changes considered as a surrogate to quantify CF. Similarly, Wu et al. (2022) employs layer-wise and task-wise probing to analyze CF in each layer for previously learned tasks. Luo et al. (2023b) propose using a series of probing tasks to evaluate the knowledge stored in LLMs and analyze the generality of the models. However, their study is limited to classification tasks and encoder-only model architectures. To the best of our knowledge, we are the first to evaluate the forgetting of general knowledge in generative large language models during continual instruction tuning.
@@ 3 Method
Set Elements
DK STEM, Social, Human, Other
Rs BoolQ, PIQA, Winogrande, Hellaswag, MathQA, Mutual
RC RACE-high, RACE-middle
Socioeconomic, Disability, Age
@@ 3.1 Continual Tasks
Following Scialom et al. (2022) (Scialom et al., 2022), we adopt instruction tasks dissimilar to the training and evaluation tasks of BLOOMZ and mT0. Specifically, we select 5 tasks from Scialom et al. (2022) as follows:
@@
During the instruction tuning, we first add a general prompt template to the beginning of the data: ‘Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request…’ followed by a specific prompt for each task. We adopt the specific prompts designed by Scialom et al. (2022) (Scialom et al., 2022) and 100,000 data samples are used for training. The details of the instruction are shown in Appendix A. For simplicity, we train the model on one instruction task order for the empirical study: Simp $\to$ Emdg $\to$ InqQG $\to$ Exp $\to$ HGen.
@@ 3.3 Evaluation Metric
We define the forgetting metric $FG$, which is the average decrease of $R^{e}_{m}$, as a surrogate metric to evaluate the forgetting:
$\centering FG_{i}=\frac{1}{|E_{i}|}\sum_{e\in E_{i}}\frac{1}{N}\sum_{m=1}^{N}\frac{R^{e}_{o}-R^{e}_{m}}{R^{e}_{o}}*100\%,\@add@centering$ (1)
@@
For most multi-choice evaluation elements, we adopt the accuracy in the zero-shot setting to measure the model performance, including MathQA, Hellaswag, BoolQ, PIQA, Mutual, Winograde, and RACE. For MMLU, we adopt the 5-shot setting for evaluation. For CrowsPairs, we follow Nangia et al. (2022) (Nangia et al., 2020) to measure the model preference for the stereotypical sentence based on the perplexity of the given stereotypical and anti-stereotypical sentences, where a larger value means a stronger bias in the language.
@@ 4.1 Large Language Models
BLOOMZ (Muennighoff et al., 2022) is a decoder-only model fine-tuned on multilingual tasks with English prompts based on the model BLOOM. The model scales range from 560M to 176b, providing a test bed for analyzing the forgetting phenomenon across different scales. Specifically, in this study, we continually train BLOOMZ on the scale 1.1b, 1.7b, 3b, and 7.1b  [^3] https://huggingface.co/bigscience/bloomz because of the limitation of computation resources.
@@ 4.2 Implementation
We train our model on 8 GPU (Tesla A100 40G) using the Adam optimizer (Kingma & Ba, 2014) (the models in 1b level are trained on 4 GPU for saving resources). For all the models, the batch size is 4 on each device, the learning rate is 2e-5, and the scheduler is set constant for BLOOMZ and mT0 following (Muennighoff et al., 2022). In LLAMA and ALPACA, we follow the hyperparameter of Taori et al. (2023) (Taori et al., 2023) that the scheduler is cosine and learning rate is 2e-5.  [^7] Note that these hype-parameters are set according to the initial work, but may still introduce some effects in the fine-tuned performance. The max sequence length of the inputs is 512. We train our model 3 epochs and the final checkpoints are used for evaluation.
@@ 5 Experimental Results and Analysis
Simp (SARI) Emdg (BS) InqQG (BS) Exp (BS) HGen (R1)
mT0-3.7b 39.01 39.92 48.29 51.70 52.66 56.25 51.62 61.91 30.50 31.77
BLOOMZ-3b 37.95 46.72 46.10 53.27 49.06 59.72 49.75 67.10 27.88 31.72
LLAMA-7b 43.02 46.92 44.28 49.54 43.90 47.54 50.72 54.22 32.42 33.80
ALPACA-7b 45.37 48.22 52.56 54.70 56.91 62.13 52.66 70.49 32.06 36.73
@@
5.1 Main Results
Firstly, we show the results of the instruction tuning during continual instruction tuning (Table 3). We mainly present the initial performance and the performance after tuning on the corresponding task to demonstrate the effectiveness of instruction tuning. We use the metrics to measure the model performance on each task following Scialom et al. (2022) (Scialom et al., 2022). In particular, we use SARI for Simp, BLUEScore for Emdg, InqQA, Exp, and Rouge-1 for HGen. For example, the initial BLOOMZ-7.1b achieves 51.47% BLUEScore in the Exp task, but after continual tuning the model on Simp, Emdg, InqQG, and Exp, the model achieves 68.71% on the Exp task. These improvements demonstrate that the model can benefit from the instruction tuning process and achieve significantly better performance on the instruction tasks.
Table 5: FG values for Bias in LLMs during continual instruction tuning. $R^{s}_{0}$ and $R^{s}_{-1}$ refer to the evaluation results at the beginning and the end of instruction tuning.
$R^{s}_{0}$ $R^{s}_{-1}$ $FG$
@@ 5.1 Main Results
Next, Figure 3 displays the FG values of BLOOMZ-1.1b and BLOOMZ-7.1B. As we can observe, the performance gradually decreases as we continually tune the model with instruction tasks. For instance, the performance of BLOOMZ-7.1b on MMLU-SocialScience in Figure 3 drops from 36.18% to 26.06% after continual training. The declining performance in LLMs indicates the presence of the catastrophic forgetting (CF) problem during the continual instruction tuning process. Moreover, as more instruction tasks are introduced, the general knowledge suffers more significant forgetting. We can also notice that the performance of the BLOOMZ-7.1b model drops more drastically in these evaluation tasks, as we adopt the same y-axis scale for both BLOOMZ-1.1b and BLOOMZ-7.1b in Figure 3. For example, the performance on MMLU-Other drops from 36.18% to 26.35% in BLOOMZ-7.1b, while it drops from 30.58% to 25.97% in BLOOMZ-1.1b after continual training. Although the initial performance of BLOOMZ-7.1b is more substantial compared to that of BLOOMZ-1.1b, the performance of both models ends at relatively similar values, which would result in a more significant FG value for BLOOMZ-7.1B.
Figure 4: The detailed results of knowledge evolution between BLOOMZ and mT0 in the comparable model scale. The first row refers to the model BLOOMZ-3b and the second refers to mT0-3,7b. The first to third columns are the FG values of domain knowledge (MMLU), reasoning and reading comprehension, respectively.
5.2 Effect of Scales
We visualize the FG values of domain knowledge, reasoning, and reading comprehension with respect to model scales in Figure LABEL:DK. We can observe that the forgetting phenomenon becomes increasingly severe as the model scale increases. For example, the FG values in domain knowledge are 9.54%, 10.72%, 14.63%, and 18.37% in BLOOMZ-1.1b, 1.7b, 3b, and 7.1b, respectively. BLOOMZ-7.1b suffers the most drastic forgetting. As shown in Table 1, the initial performance $R_{o}^{s}$ is boosted by the increasing model scale, but the final performance is relatively similar across different scales, which may explain the varying extent of forgetting. For example the task in domain knowledge are mostly multi-choice problems with 4 options and all the model tend to achieve a nearly random guess performance, but BLLOMZ-7b perform much significant initially. It may imply that these models still shift parameters in a large extent to fit the instruction tasks. The same pattern can also be observed in the mT0-1.2b and mT0-3.7b models, as shown in Table 4. Regarding the bias in LLMs, FG values do not correlate with the model scales, which is also reflected in the initial model performance. In other words, there is no evident correlation between the initial performance $R_{o}^{s}$ in bias and the model scales. This finding suggests that the degree of bias in LLMs is not directly related to their size, and increasing the model scale does not necessarily lead to a corresponding increase or decrease in bias.
@@ 5.3 Effect of Model Architecture
We also compare the forgetting phenomenon of different model architectures in Figure 4. As observed, at comparable model scale, BLOOMZ-1.1b and mT0-1.2b achieve similar FG values in domain knowledge, reasoning, and reading comprehension. However, as the scale increases to 3b, BLOOMZ-3b suffers less forgetting compared to mT0-3.7B. For example, the FG value of BLOOMZ-3b is 11.09 which is 5.64 lower than that of mT0-3.7b. These results suggest that BLOOMZ, which has a decoder-only model architecture, can maintain more knowledge during continual instruction tuning. This difference may be attributed to the autoregressive nature of the model or the differences in training objectives. Furthermore, the results imply that as the model scale increases, decoder-only models may suffer from less catastrophic forgetting compared to encoder-decoder models. As we observe, the knowledge degraded more drastically in mT0.